# The Name of the Title Is Hope

### v2

@@ Abstract.
Continual learning(CL) is a useful technique to acquire dynamic knowledge continually. Although powerful cloud platforms can fully exert the ability of CL,e.g., customized recommendation systems, similar personalized requirements for edge devices are almost disregarded. This phenomenon stems from the huge resource overhead involved in training neural networks and overcoming the forgetting problem of CL. This paper focuses on these scenarios and proposes a compact algorithm called LightCL. Different from other CL  methods bringing huge resource consumption to acquire generalizability among all tasks for delaying forgetting, LightCL  compress the resource consumption of already generalized components in neural networks and uses a few extra resources to improve memory in other parts. We first propose two new metrics of learning plasticity and memory stability to seek generalizability during CL. Based on the discovery that lower and middle layers have more generalizability and deeper layers are opposite, we Maintain Generalizability by freezing the lower and middle layers. Then, we Memorize Feature Patterns to stabilize the feature extracting patterns of previous tasks to improve generalizability in deeper layers. In the experimental comparison, LightCL  outperforms other SOTA methods in delaying forgetting and reduces at most 6.16$\times$ memory footprint, proving the excellent performance of LightCL  in efficiency. We also evaluate the efficiency of our method on an edge device, the Jetson Nano, which further proves our method’s practical effectiveness.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXX11footnotetext: Equal contribution22footnotetext: Correspondence to chengfei@xidian.edu.cn
@@ 1. Introduction
The wave of deep learning has significantly boosted the deployment of the neural network from cloud platforms to edge devices(Lin et al., 2023, 2022). As the dynamic realistic environment requires the adaptive ability of models, a potential technique known as continual learning (CL) empowers these models to continually acquire, update, and accumulate knowledge on the device from the data by the training behavior. Although powerful cloud platforms can leverage CL  to adapt to service expansion, e.g., personalized recommendation systems(Lee et al., 2020), the similar and intense personalized requirements for edge devices are almost disregarded by academia and industry. This overlook stems from scarce resources in edge devices, such as limited computation, constrained storage, and insufficient memory, which significantly hinders the further application of CL. Hence, efficient CL  on edge devices becomes a considerable challenge.
@@
Most existing CL  methods ignore edge scenarios and only concentrate on addressing the catastrophic forgetting problem(McCloskey and Cohen, 1989; Kirkpatrick et al., 2017) – model training on a new task tends to ”forget” the knowledge learned from previous tasks. Specifically, these methods explore two factors of generalizability(Wang et al., 2023), learning plasticity and memory stability, and make a trade-off between them to delay forgetting. Note that learning plasticity denotes the adaptation to the new task, while memory stability denotes the memory capacity of previous tasks. Hence, most of CL  methods(Buzzega et al., 2020; Chaudhry et al., 2019; Benjamin et al., 2019; Kirkpatrick et al., 2017; Rusu et al., 2016) simultaneously train on the trace of historical information and new data for generalizability. It is well known that the training overhead of the neural network is already significant. Meanwhile, learning from history in CL  requires additional resources for the training process. Thus, the casual CL methods are unsuitable for resource-intensive scenarios. To mitigate this problem, compression techniques are inevitably introduced to the on-device CL. Especially, sparse learning(Han et al., 2015) has become a popular paradigm of the efficient CL(Wang et al., 2022; Ren and Honavar, [n. d.]). Although they(Wang et al., 2022; Ren and Honavar, [n. d.]) greatly reduce computation by the sparsity, intermediate feature maps cannot be effectively reduced, leading to an enormous memory footprint during training.
@@
Let’s dive into the challenges of the on-device CL. Firstly, the forgetting problem is the first and foremost issue in CL, whether on edge devices or cloud platforms. Most CL methods train on historical information and new data together for generalizability to delay forgetting. This training behavior requires more significant resources than usual, including intermediate feature maps and gradients in memory, as well as additional computations for gradient calculations. This greatly hinders the efficiency. Secondly, the efficiency of CL heavily influences its prevalence in edge devices. The computation and memory footprint are two main factors that hinder the deployment. Meanwhile, memory rather than computation has become the primary bottleneck in AI applications, and the time taken for memory transfers greatly limits computational capacity(Gholami et al., 2024), which impedes the process of overcoming the forgetting during CL. Actually, the above two challenges are not orthogonal. The former inevitably interferes with the latter, and vice versa. This tangle prompts us to solve two challenges together. Considering the neural network has the potential to train in new tasks with the help of previous knowledge during CL, it absolutely has redundancy of generalizability. In other words, if we could evaluate the distribution of generalizability in the neural network during CL, we can compress the training redundancy with the over-strong generalizability and use just a few extra resources to maintain memory in less-generalized parts, which hopefully gives us a more favorable and efficient CL on edge devices.
@@
Therefore, we start with the two factors of generalizability, learning plasticity and memory stability, and design their metrics to quantitatively assess the generalizability of neural networks during CL. This evaluation demonstrates that the generalizability of different layers in a neural network exhibits significant variation. In particular, lower and middle layers have more generalizability, and deeper layers are the opposite. According to this conclusion, we Maintain Generalizability by easily freezing lower layers to delay forgetting and greatly reduce memory footprint and computation. Furthermore, for less-generalized parts in deeper layers, we regulate their feature maps, necessitating fewer resources, to Memorize Feature Patterns of previous tasks for the promotion of generalizability. The resource consumption of Memorize Feature Patterns is much smaller than its reduction of Maintain Generalizability. Thus, our method can achieve great performance with limited resources. Since our method achieves high efficiency only by exploring the network structure in CL, it can be easily combined with other compression techniques. As shown in Figure  1, the performance of LightCL outperforms the state-of-the-art efficient CL method SparCL and reduces at most $6.16$$\times$ memory footprint.
@@
  • It is the first study to propose two new metrics of learning plasticity and memory stability to quantitatively evaluate generalizability for efficient CL. We conclude that lower and middle layers have more generalizability, and deeper layers are the opposite.
  • Based on the evaluation, LightCL  efficiently overcomes catastrophic forgetting through Maintaining Generalizability and Memorize Feature Patterns.
@@ 2. RELATED WORK
Continual Learning Methods: CL  methods can be categorized into replay methods(Buzzega et al., 2020; Chaudhry et al., 2019; Benjamin et al., 2019; Hayes et al., 2020; Cha et al., 2021), regularization-based methods(Kirkpatrick et al., 2017; Li and Hoiem, 2018; Zenke et al., 2017; Aljundi et al., 2018), and architecture-based methods(Mallya and Lazebnik, 2018; Rusu et al., 2016; Wang et al., 2020). Replay methods address catastrophic forgetting by replaying a subset of old data stored in storage. They must continually store samples when they meet new tasks, and bring extra computation and memory footprint for training on old data. Regularization-based methods are designed to restrict the updates of neural networks through regularization, preserving valuable information about previous tasks. Most of them(Kirkpatrick et al., 2017; Zenke et al., 2017; Aljundi et al., 2018) limit parameter changes necessary for old tasks, leading to extra consumption for recording the importance of every parameter. Architecture-based methods can prevent forgetting by distributing different model structures to each task. Similar work(Rusu et al., 2016) instantiates a new network for each new task, while others(Wang et al., 2020) divides a network into task-specific parts through pruning. Hence, most of them grow the model size continually or train on a whole dense model to learn new tasks, which causes significant resource overhead during the CL  process. Meanwhile, they must know the identity of the task before the model’s evaluation, which is unsuitable for edge scenarios.
Efficient Learning: Many efficient algorithms aim at the deployment of neural networks on edge devices, such as pruning(Han et al., 2016; Li et al., 2017), quantization(Gong et al., 2014) and etc. Pruning compresses networks by discarding weight redundancies, while quantization compresses the bit-width of weights. These methods mainly accelerate the inference procedure on edge(Shuvo et al., 2023), and degrade the ability of compressed models to accommodate dynamic environments. That’s because large training consumption hinders deployment development, such as massive usage of intermediate feature maps and gradients to memory footprint and more training FLOPs on computation. Although (Lin et al., 2022) can train in resource-limited settings, they focus more on training a single task and don’t consider catastrophic forgetting between multiple tasks. To cope with catastrophic forgetting of CL  in the edge device scenario, SparCL(Wang et al., 2022), combined with the dynamic pruning method, greatly reduces training FLOPs and delays the forgetting problem. However, SparCL still occupies massive memory footprints for intermediate feature maps.
@@ 3. CL Setting
In the CL  setting, tasks $\{1,2,...,T\}$ are executed sequentially during training, and data from previous tasks are inaccessible when training on the new task. Specifically, the t-th task represents $\mathcal{D}_{t}=\{(x_{t}^{i},y_{t}^{i})\}^{n_{t}}_{i=1}$, where $n_{t}$ is the sample number of t-th task and $(x_{t}^{i},y_{t}^{i})$ is input-label pair. $\theta_{i}^{t}$ represents the i-th layer of the neural network after training (or during) t-th task. The goal of CL  is to let the AI model perform well on all tasks when learning a new task. Overall performance is typically evaluated by average accuracy (AA)(Wang et al., 2023, 2022) on all tests after training. AA defined as $AA=\frac{1}{T}\sum_{i=1}^{T}a_{i}(\theta^{T})$, where $a_{i}(\theta)$ represents the testing accuracy of model $\theta$ on i-th task. There are two main CL  scenarios(Wang et al., 2022): 1) Task incremental learning (TIL), the identity $t$ of the task $\mathcal{D}_{t}$ are available during training and testing; 2) Class incremental learning (CIL), the task identity $t$ are only available during training. This paper evaluates CL  methods under TIL and CIL scenarios.
@@ 4. Analysis of Generalizability
In this section, we analyze the distribution of generalizability in the neural network during CL. Training consumption on generalized components in the neural network can be a kind of redundancy since they already have the ability to fit new tasks. That’s the motivation for evaluating generalizability. Considering generalizability means can be utilized for both previous and new tasks, we utilize memory stability (MS) and learning plasticity (LP) to represent these two different characters respectively. Similar work(Ramasesh et al., 2021) analyzes the forgetting parts in the neural network during CL. In fact, forgetting means strong learning plasticity and less memory stability, so generalizability and forgetting are not orthogonal and we cannot infer generalizability from it(Ramasesh et al., 2021). Thus, we need to design two metrics of LP and MS to quantitatively evaluate generalizability.
MS measures the loss of previous knowledge and LP measures adaptation to new knowledge. We evaluate them from the perspective of layers in the neural network. Specifically, we replace the parameters of i-th layer in the model $\theta^{T}$ with the parameters from the previous model $\theta^{T-1}$. The new model $\theta^{new}$ can be described as
(1) $\begin{gathered}\theta^{new}=\begin{cases}\theta^{new}_{j}=\theta^{T}_{j}&\text{if }j\neq i\\ \theta^{new}_{i}=\theta^{T-1}_{i}&\text{otherwise.}\end{cases}\end{gathered}$
Then, we separately test the accuracy of $\theta^{new}$ on (T-1)-th task and T-th task. Two metrics of LP and MS can be described as
(2) $\begin{gathered}\begin{cases}MS=a_{T-1}(\theta^{new})-a_{T-1}(\theta^{T})\\ LP=a_{T}(\theta^{new})-a_{T}(\theta^{T}).\end{cases}\end{gathered}$
If the value of MS is high, it indicates that the previous knowledge in i-th layer has been lost to a large extent, suggesting that the memory stability of i-th layer is weak. The low value of LP indicates that the previous i-th layer doesn’t have a good adaptation for the new tasks, meaning that the previous i-th layer has weak plasticity.
@@
For more reasonable, we evaluate the above metrics using different class numbers for multiple tasks across $5$ runs on two typical backbones, ResNet-18(He et al., 2016) and VGG16(Simonyan and Zisserman, 2015). As shown in Figure  2, different numbers of classes exhibit a similar phenomenon during CL learning. Lower layers have small values of MS and high values of LP. In other words, the lower layers have high values of memory stability and learning plasticity, which exhibit good generalizability. The middle layers in ResNet-18 display good generalization ability, while the counterpart of VGG16 exhibits poor memory stability (second row in Figure 2(2(a))). It should be noted that LP suggests that middle layers already exhibit learning plasticity before being trained on the new task in Figure 2(2(b)), indicating that learning the new task will impair the stability of the middle layers and degrade the existing generalizability. This phenomenon proves the opinion that forgetting parts is not orthogonal to generalized parts. For the deeper layers, the values of MS are higher and the values of LP are lower. This suggests that the deeper layers are less-generalizability and forgetting parts. From the above analysis, we conclude that during CL, lower and middle layers have strong generalizability, and deeper layers have less generalizability. This conclusion guides the design of LightCL  in Section 5.
5. Methodology

### v3

@@ Abstract.
Continual learning (CL) is a technique that enables neural networks to constantly adapt to their dynamic surroundings. Despite being overlooked for a long time, this technology can considerably address the customized needs of users in edge devices. Actually, most CL  methods require huge resource consumption by the training behavior to acquire generalizability among all tasks for delaying forgetting regardless of edge scenarios. Therefore, this paper proposes a compact algorithm called LightCL, which evaluates and compresses the redundancy of already generalized components in structures of the neural network. Specifically, we consider two factors of generalizability, learning plasticity and memory stability, and design metrics of both to quantitatively assess generalizability of neural networks during CL. This evaluation shows that generalizability of different layers in a neural network exhibits a significant variation. Thus, we Maintain Generalizability by freezing generalized parts without the resource-intensive training process and Memorize Feature Patterns by stabilizing feature extracting of previous tasks to enhance generalizability for less-generalized parts with a little extra memory, which is far less than the reduction by freezing. Experiments illustrate that LightCL  outperforms other state-of-the-art methods and reduces at most $6.16$$\times$ memory footprint. We also verify the effectiveness of LightCL  on the edge device.
journalyear: 2025copyright: acmlicensedconference: 30th Asia and South Pacific Design Automation Conference; January 20–23, 2025; Tokyo, Japanbooktitle: 30th Asia and South Pacific Design Automation Conference (ASPDAC ’25), January 20–23, 2025, Tokyo, Japandoi: 10.1145/3658617.3697784isbn: 979-8-4007-0635-6/25/0111footnotetext: Equal contribution22footnotetext: Correspondence to chengfei@xidian.edu.cnfootnotetext: Our code is available at: https://github.com/INV-WZQ/LightCL
@@ 1. Introduction
The wave of deep learning has significantly boosted the deployment of neural networks from cloud platforms to edge devices (Lin et al., 2023, 2022). As the dynamic realistic environment requires the adaptive ability of models, a potential technique known as continual learning (CL) empowers these models to continually acquire, update, and accumulate knowledge on the device from the data by the training behavior. Although powerful cloud platforms can leverage CL  to adapt to service expansion, e.g., personalized recommendation systems (Lee et al., 2020), the similar and intense personalized requirements for edge devices are almost disregarded by academia and industry. This overlook stems from scarce resources in edge devices, such as limited computation, constrained storage, and insufficient memory, which significantly hinders the further application of CL. Hence, the efficient CL on the edge device becomes a considerable challenge.
@@
Most existing CL  methods ignore edge scenarios and only concentrate on addressing the catastrophic forgetting problem (McCloskey and Cohen, 1989; Kirkpatrick et al., 2017) – model training on a new task tends to ”forget” the knowledge learned from previous tasks. Specifically, these methods explore two factors of generalizability (Wang et al., 2023), learning plasticity and memory stability, and make a trade-off between them to delay forgetting. Note that learning plasticity denotes the adaptating ability to the new task, while memory stability denotes the memory capacity of previous tasks. Unfortunately, most of CL  methods (Buzzega et al., 2020; Chaudhry et al., 2019; Benjamin et al., 2019; Kirkpatrick et al., 2017; Rusu et al., 2016) simultaneously train on the trace of historical information and new data for generalizability. It is well known that the training overhead of the neural network is already significant. Meanwhile, learning from history in CL  requires additional resources for the training process. Accordingly, the casual CL methods are unsuitable for resource-intensive scenarios. To mitigate this problem, compression techniques are inevitably introduced to the on-device CL. Especially, sparse learning (Han et al., 2015) has become a popular paradigm of the efficient CL  (Wang et al., 2022; Ren and Honavar, 2024). Although they (Wang et al., 2022; Ren and Honavar, 2024) greatly reduce computation by removing unimportant weights, intermediate feature maps cannot be effectively reduced, leading to an enormous consumption of memory footprint during training.
@@
Before introducing our work, we will give a comprehensive analysis of the challenges encountered. Firstly, the forgetting problem is the first and foremost issue of CL, whether on edge devices or cloud platforms. Most CL methods train on historical information and new data together for generalizability to delay forgetting. This training behavior requires more significant resources than usual, including more intermediate feature maps and gradients in memory, as well as additional computations for gradient calculations. This greatly hinders the efficiency. Secondly, the efficiency of CL heavily influences its prevalence in edge devices. The computation and memory footprint are two main factors that hinder the deployment. Meanwhile, memory rather than computation has become the primary bottleneck in AI applications, and the time taken for memory transfers greatly limits computational capacity (Gholami et al., 2024), which impedes the process of overcoming forgetting during CL.
Actually, the above two challenges are not orthogonal. The former inevitably interferes with the latter, and vice versa. This tangle prompts us to solve two challenges together. Considering the neural network has the potential to train in new tasks with the help of previous knowledge during CL, it absolutely has the redundancy of generalizability. In other words, if we could evaluate the distribution of generalizability in the neural network during CL, we can compress the training redundancy with the over-strong generalizability and use just a few extra resources to maintain memory in less-generalized parts, which hopefully gives us a more favorable and efficient CL method on edge devices.
@@
Therefore, we start with the two factors of generalizability, learning plasticity and memory stability, and design their metrics to quantitatively assess the generalizability of neural networks during CL. This evaluation demonstrates that the generalizability of different layers in a neural network exhibits a significant variation. In particular, lower and middle layers have more generalizability, and deeper layers are the opposite. According to this conclusion, we Maintain Generalizability by easily freezing lower layers to delay forgetting problem and greatly minimize the consumption of memory footprint and computation. Furthermore, as for less-generalized parts in deeper layers, we regulate their feature maps, necessitating fewer resources than weights, to Memorize Feature Patterns of previous tasks for the promotion of generalizability. The resource consumption of Memorize Feature Patterns is significantly less than its reduction of Maintain Generalizability. Thus, our method can achieve great performance with limited resources. Since our method achieves high efficiency only by exploring the network structure in CL, it can be easily integrated with other compression techniques. As shown in Figure  1, with the same sparsity ratio, the performance of LightCL outperforms the state-of-the-art efficient CL method SparCL and reduces at most $6.16$$\times$ memory footprint.
@@
  • It is the first study to propose two new metrics of learning plasticity and memory stability to quantitatively evaluate generalizability for efficient CL. We conclude that lower and middle layers have more generalizability, while deeper layers hold the opposite.
  • Based on the evaluation, LightCL  efficiently overcomes catastrophic forgetting through Maintain Generalizability and Memorize Feature Patterns.
@@ 2. RELATED WORK
Continual Learning Methods. CL  methods can be categorized into replay methods (Buzzega et al., 2020; Chaudhry et al., 2019; Benjamin et al., 2019; Hayes et al., 2020; Cha et al., 2021), regularization-based methods (Kirkpatrick et al., 2017; Li and Hoiem, 2018; Zenke et al., 2017; Aljundi et al., 2018), and architecture-based methods (Mallya and Lazebnik, 2018; Rusu et al., 2016; Wang et al., 2020). Replay methods address catastrophic forgetting by replaying a subset of old data stored in storage. They must continually store samples when they meet new tasks, and bring extra computation and memory footprint for training on old data. Regularization-based methods are designed to restrict the updates of neural networks through regularization, preserving valuable information about previous tasks. Most of them (Kirkpatrick et al., 2017; Zenke et al., 2017; Aljundi et al., 2018) limit changes of parameters that are important to old tasks. This manner leads to extra consumption for recording the importance of every parameter. Architecture-based methods can mitigate forgetting by distributing different model structures to each task. Similar work (Rusu et al., 2016) instantiates a new network for each new task, while others (Wang et al., 2020) divide a network into task-specific parts through pruning. Hence, most of them grow the model size continually or train on a whole dense model to fit new tasks, which causes significant resource overhead during the CL  process. Meanwhile, they must acquire the identity of the task before an evaluation, which is unsuitable for edge scenarios.
Efficient Learning. Many efficient algorithms aim at the deployment of neural networks on edge devices, such as pruning (Han et al., 2016; Li et al., 2017), quantization (Gong et al., 2014), etc. Pruning compresses networks by discarding redundancies of weights, while quantization reduces the bit-width of weights. These methods mainly accelerate the inference procedure on edge (Shuvo et al., 2023), and degrade the ability of compressed models to accommodate dynamic environments. This is because a large training burden hinders deployment development, such as the extensive use of intermediate feature maps and gradients, which increases memory footprint and requires more training FLOPs for computation. Although (Lin et al., 2022) can train in resource-limited settings, they focus more on training a single task and do not consider catastrophic forgetting between multiple tasks. To cope with catastrophic forgetting of CL  in the edge device scenario, SparCL (Wang et al., 2022), combined with the dynamic pruning method, greatly reduces training FLOPs and delays the forgetting problem. However, SparCL still occupies massive memory footprints for intermediate feature maps.
@@ 3. CL Setting
In the CL  setting, tasks $\{1,2,...,T\}$ are executed sequentially during training, and data from previous tasks are inaccessible when training on the new task. Specifically, the t-th task represents $\mathcal{D}_{t}=\{(x_{t}^{i},y_{t}^{i})\}^{n_{t}}_{i=1}$, where $n_{t}$ is the number of samples of t-th task and $(x_{t}^{i},y_{t}^{i})$ is input-label pair. $\theta_{i}^{t}$ represents the i-th layer of the neural network after (or during) training t-th task. The goal of CL  is to let the AI model perform well on all tasks when learning a new task. Overall performance is typically evaluated by average accuracy (AA) (Wang et al., 2023, 2022) on all tests after training. AA is defined as $AA=\frac{1}{T}\sum_{i=1}^{T}a_{i}(\theta^{T})$, where $a_{i}(\theta)$ represents the testing accuracy of model $\theta$ on i-th task. There are two main CL  scenarios (Wang et al., 2022): 1) Task incremental learning (TIL), the identity $t$ of the task $\mathcal{D}_{t}$ is available during training and testing; 2) Class incremental learning (CIL), the task identity $t$ are only available during training. This paper evaluates CL  methods under both TIL and CIL scenarios.
@@ 4. Analysis of Generalizability
In this section, we analyze the distribution of generalizability for the neural network during CL. Training consumption on generalized components in the neural network can be a kind of redundancy since certain generalized components contain previous knowledge while already have the ability to fit new tasks prior to encountering them. This is the motivation for evaluating generalizability. As (Wang et al., 2023) points out that generalizability can be utilized for both previous and new tasks during CL, memory stability (MS) and learning plasticity (LP) are two different characteristics of it. MS denotes the loss of previous knowledge, and LP denotes the adaptation to new knowledge. Therefore, we need to analyze the generalizability from these two perspectives. Although similar work (Ramasesh et al., 2021) makes a deep analysis in continual learning, it focuses more on analyzing forgetting parts in the neural network. In fact, forgetting means strong learning plasticity and weak memory stability, so generalizability and forgetting are not orthogonal and we cannot infer generalizability from the previous work (Ramasesh et al., 2021). Hence, we need to design two metrics of LP and MS to quantitatively evaluate generalizability.
Intuitively, to analyze MS and LP in specific structures of neural networks, the possible way is to analyze the effectiveness of specific structures on previous and new tasks. Through testing on previous tasks, we are able to evaluate the loss of previous knowledge in certain structures. By testing on new tasks, we can also evaluate their adapting ability to new knowledge. Specifically, we quantitatively evaluate the above metrics of each layer of the neural network. We replace the parameters of i-th layer in the model $\theta^{T}$ with the parameters of the same layer from the previous model $\theta^{T-1}$. The new model $\theta^{new_{i}}$ can be described as
(1) $\begin{gathered}\theta^{new_{i}}=\begin{cases}\theta^{new_{i}}_{j}=\theta^{T}_{j}&\text{if }j\neq i,\\ \theta^{new_{i}}_{i}=\theta^{T-1}_{i}&\text{otherwise.}\end{cases}\end{gathered}$
Then, we separately test the accuracy of $\theta^{new_{i}}$ on (T-1)-th task and T-th task. Two metrics of LP and MS can be described as
(2) $\begin{gathered}\begin{cases}MS=a_{T-1}(\theta^{new_{i}})-a_{T-1}(\theta^{T}),\\ LP=a_{T}(\theta^{new_{i}})-a_{T}(\theta^{T}).\end{cases}\end{gathered}$
If the value of MS is high, it indicates that the previous knowledge in i-th layer has been lost to a large extent, suggesting that the memory stability of i-th layer is weak. The low value of LP indicates that the previous i-th layer does not have a good adaptation for the new tasks, meaning that the previous i-th layer has weak plasticity.
@@
In order to make our experiment more reasonable, we evaluate the above metrics using different class numbers for multiple tasks across $5$ runs with two typical backbones, ResNet-18 and VGG16 (Simonyan and Zisserman, 2015). As shown in Figure  2, different numbers of classes exhibit a similar phenomenon that, in lower layers, the values of MS are small while the values of LP are high. In other words, the lower layers have strong memory stability and learning plasticity, which exhibit strong generalizability. The middle layers in ResNet-18 display strong generalization ability, while the counterpart of VGG16 exhibits poor memory stability, shown in the second row of Figure 2(2(a)). As shown in Figure 2(2(b)), high value in LP of middle layers is important to note that middle layers already exhibit learning plasticity before being trained on the new task, indicating that learning the new task will impair the stability of the middle layers and degrade the existing generalizability. This phenomenon also proves the opinion that forgetting parts is not orthogonal to generalized parts since they may display different features during learning process like middle layers. For the deeper layers, the values of MS are higher and the values of LP are lower. This suggests that the deeper layers are less-generalizability and forgetting parts. From the above analysis, we conclude that during CL, lower and middle layers have strong generalizability, and deeper layers have less generalizability. This conclusion guides the design of LightCL  in Section 5.
5. Proposed Architecture